Steinberg Corporation and Dietrich Corporation are identical firms except that Dietrich is more levered. Both companies will remain in business for one more year. The companies' economists agree that the probability of the continuation of the current expansion is 70 percent for the next year, and the probability of a recession is 30 percent. If the expansion continues, each firm will generate earnings before interest and taxes (EBIT) of $3.8 million. If a recession occurs, each firm will generate earnings before interest and taxes (EBIT) of $1.2 million. Steinberg's debt obligation requires the firm to pay $920,000 at the end of the year. Dietrich's debt obligation requires the firm to pay $1.3 million at the end of the year. Neither firm pays taxes. Assume a discount rate of 15 percent. a-1. What are the current market values of Steinberg's equity and debt? (Enter your answers in dollars, not millions of dollars. Do not round intermediate calculations and round your answers to the nearest whole dollar, e.g., 1,234,567.) Steinberg Equity value $ Debt value $ a-2. What are the current market values of Dietrich's equity and debt? (Enter your answers in dollars, not millions of dollars. Do not round intermediate calculations and round your answers to the nearest whole dollar, e.g., 1,234,567.) Dietrich Equity value $ Debt value $ b. Steinberg’s CEO recently stated that Steinberg’s value should be higher than Dietrich’s since the firm has less debt, and, therefore, less bankruptcy risk. Do you agree or disagree with this statement? Disagree Agree
Answer #1

a-1)

Steinberg
Particulars Expansion Recession
EBIT $3,800,000 $1,200,000
Less: Earnings to Bondholders $920,000 $920,000
Earnings available to equityholders $2,880,000 $280,000

Market value of debt = [ $920,000 x 0.70 + $920,000 x 0.30 ] / (1 + 0.15) = $800,000

Market value of equity = [ $2,880,000 x 0.70 + $280,000 x 0.30 ] / (1 + 0.15) = $1,826,087

a-2)

Dietrich
Particulars Expansion Recession
EBIT $3,800,000 $1,200,000
Less: Earnings to Bondholders $1,300,000 $1,200,000 (cannot pay more than earnings)
Earnings available to equityholders $2,500,000 $0

Market value of debt = [ $1,300,000 x 0.70 + $1,200,000 x 0.30 ] / (1 + 0.15) = $1,104,348

Market value of equity = [ $2,500,000 x 0.70 + $0 x 0.30 ] / (1 + 0.15) = $1,521,739

c) Total value of Steinberg = $800,000 + $1,826,087 = $2,626,087

Total value of Dietrich = $1,104,348 + $1,521,739 = $2,626,087

We market values are same for both the companies, therefore, we disagree with the statement.